OpenStudy (jackie4uxd):

Calculus :Find the number of unit x that produces a maxima revenue R. R=228x^2 - 0.08x^3 x= ? units

OpenStudy (anonymous):

Are you taking Calculus?

OpenStudy (jackie4uxd):

yes

OpenStudy (anonymous):

OK, for Calculus questions, I suggest you mark your questions as Calculus and you get a better response from the experts. I should let them answer Calculus questions, but I believe all you need to do it the following. Take the first derivative of your equation. You will get a quadratic equation which you can solve for x using the quadratic equation formula. This will get you x values for the maximum and minimum values that you are seeking. Plug the x value back into your original equation to find the y value or R value in your case. Also, I suggest you use the equation editor by immediately replying to your question and clicking on the equation button.

OpenStudy (anonymous):

You can also ping one of the experts. Unfortunately, I don't know how to do it.

OpenStudy (jackie4uxd):

Thank you ill edit my question so people know its calculus.

OpenStudy (freckles):

Hello?

OpenStudy (freckles):

Do you still need help?

OpenStudy (jackie4uxd):

hi can you help me with this problem?

OpenStudy (freckles):

I agree with marchall in taking the first derivative. Have you done that?

OpenStudy (freckles):

Have you learned power rule? Or are you just allowed to use the definition of derivative to find the derivative?

OpenStudy (jackie4uxd):

i got \[456x-0.24x^2\]

OpenStudy (freckles):

Now you want to find the critical numbers.

OpenStudy (jackie4uxd):

i learned how to find it by using the second derivative test and finding critical numbers

OpenStudy (freckles):

That means we need find when that is zero.

OpenStudy (freckles):

You can factor your expression.

OpenStudy (freckles):

There is common factor x in both terms.

OpenStudy (jackie4uxd):

i g0t 1900=x

OpenStudy (jackie4uxd):

and x=0

OpenStudy (freckles):

Those critical numbers look great.

OpenStudy (freckles):

Now as you suggested we can determine if those numbers will give us a max or a min.

OpenStudy (freckles):

By finding second derivative.

OpenStudy (jackie4uxd):

second derivative is 456-0.48x

OpenStudy (freckles):

\[f''(x)=456-0.48x \\ f''(0)=456>0 \\ f''(1900)=-456<0\]

OpenStudy (freckles):

So when f'' >0 we have what? and when f''<0 we have?

OpenStudy (freckles):

I like to go back to a real basic graph to help me remember. f(x)=x^2 has a min and f''>0 and if f(x)=-x^2 then f has a max and f''<0

OpenStudy (jackie4uxd):

f''>0 min, f''<0 max

OpenStudy (freckles):

yes exactly

OpenStudy (jackie4uxd):

Thank you so much ! this really helped a lot!

OpenStudy (freckles):

You did all the work jackie. I think you know more than you think you do.